# Dynamically Augmented CVaR for MDPs

### v3

@@ 3. The resulted finite sequence where and represents the concave piecewise linear function
Subroutine 1 is the major step in recursive computations of value functions for
@@
This section describes the properties of value functions and for DRMDP1 defined in equations (4.5), where or It also describes the properties of the sets of optimal actions for the DM. We recall that the function is concave and continuous in because the function is continuous in and because Assumption 5.1 holds. We observe that, if then the function is linear in
@@
Proof.
Proof of Theorem 5.2 For the given time horizon or and for a given sequence the algorithm sequentially generates a finite or infinite sequence of optimal actions. Let be the states of the system and be the tail risk levels, and the values are not observed by the DM when In addition to the parameters of the model, the algorithm also uses the value functions , if and it uses the value function if
The parameter takes two values: 1 and 0. means that at the current time epoch the tail risk level is known, and indicates that it is not known. If then and this is the case If then an optimal action at epoch is chosen by the algorithm in step 2.4 after the algorithm either detects or detects that belongs to a linear interval with the slope defined in formula (5.3).
Let an optimal action be selected at an epoch by step 1 or 2.4 of the algorithm. Then
where the function is defined in (4.5) with Formula (6.5) implies
@@
If there are multiple points such that that is, the case takes place, then concavity of implies that there is a maximal interval such that the function is linear in and its slope is According to Lemma 7.2, for all the sets coincide, and for Therefore, step 2.1 can be skipped, and future calculations do not depend on the particular value So, in spite of the fact that the DM does not know the states when The algorithm either calculates values or can calculate intervals to which values belong, and optimal action sets coincides for tail risk levels on these intervals.
@@ 9 Extension to Stochastic Cost Functions
In this section, we extend the results of this paper to random one-step cost functions with finite support. Arbitrary random one-step costs can be approximated by such costs.
@@ Appendix A Results on Robust MDPs Used in this Paper
This appendix provides results on RMDPs used in this paper. In general, for robust optimization problems, some parameters of the model are unknown, and the goal is to select the best solutions for the worst possible values of unknown parameters. Starting from [19] and [23] there are numerous studies of RMDPs. For an RMDP, there is an additional parameter on which one-step costs and transition probabilities depend, and at each step Nature (sometimes called Player II) selects this parameter from a set of available parameters. The DM (sometines called Player I) chooses actions by knowing the past and current states and the past actions chosen by the DM and by Nature. Nature chooses actions by knowing this information and, in addition, by knowing the current action chosen by the DM. Such models were also studied under the names of games with perfect information (see e.g., [1, 12, 20]) and minimax control [14].
@@
An RMDP is defined by a tuple Here is the state space, and there are two players: the DM and Nature. There are two decision sets and where is the set of actions for the DM, and is the set of actions for Nature. We assume that , and are nonempty Borel subsets of Polish (complete, separable, metric) spaces. The sets where are assumed to be nonempty and finite, and is the set of actions available to the DM at states For each pair where and a set of feasible actions of Nature is defined. We assume that the sets are nonempty and compact subsets of for all and for all We also assume that is a Borel subset of and the set-valued mapping is continuous. Continuity of implies Borel measurability of According to the Arsenin-Kunugui measurable selection theorem, each of the graphs and contains at least one Borel measurable function. This fact, which is needed for the existence of policies, also follows from the Kuratowski and Ryll-Nardzewski measurable selection theorem.
@@
where is a bounded continuous function, and for the infinite horizon
@@
where, in view of Berge’s maximum theorem, each function is continuous in and therefore is a continuous function. Equations (A.1) and (A.2) imply the minimax equation
Let Then all continuous functions are uniformly bounded by where Banach’s fixed point theorem applied to a space of uniformly bounded continuous functions on implies that the minimax operator applied in (A.3) to the function has a unique fixed point which is a bounded continuous function, and as In particular
@@
minimizing the right-hand side of (A.2) and for Nature, for and for
maximizing the right-hand side of (A.1). The sets are finite and, as follows from the Berge maximum theorem, the sets are compact.
For an -horizon problem with or a pair of policies is called an equilibrium for all policies and
If a pair is an equilibrium, then is called an optimal policy for the DM. Optimal policies for Nature will be defined later in this section.
The key facts are that for a finite or infinite horizon there are Markov optimal policies for the DM and for Nature, and their optimality can be defined in a stronger sense called persistent optimality. In addition, these policies have certain structural properties, and for there are deterministic optimal policies. In order to formulate these facts as a theorem, we need to introduce additional notations and definitions.
Let the DM and Nature reset the clock to 0 at time Let the history prior to time be and let and be the histories the DM and Nature observe respectively during the following units of time after the clock is reset to 0. If the clock is not not reset, then at time the history would be for the DM and for Nature.
@@
At the initial time step 0, Nature knows the initial state and the initial action chosen by the DM. Let us define the expected total cost for an -horizon problem, where or if the DM and Nature play policies and respectively, the initial state is and the DM played an action at the initial time step,
where for all The value of does not depend on the distribution If is an equilibrium pair for the horizon then
A policy for Nature is called optimal, if there exists a policy for the DM such that is a equilibrium pair, and
In view of (A.8), a policy is optimal, if there exists a policy such that inequalities (A.9) and the right inequalities in (A.7) hold.
A policy for the DM ( for Nature) is called persistently optimal for an -horizon problem with or if, for each nonnegative integer and for each
the policy () is optimal for the horizon
@@
(i) a policy for the DM ( for Nature) is persistently optimal if and only if
() for all nonnegative integers and for all
and
(ii) there exist Markov persistently optimal policies for the DM and for Nature, and a Markov policy for the DM ( for Nature) is persistently optimal if and only if () for all nonnegative integers and for all and
(iii) if and are optimal policies for the DM and Nature respectively, then
(iv) for there exist deterministic optimal policies for the DM and for Nature, and a deterministic policy for the DM ( for Nature) is optimal if and only if () for all
@@
The proof is based on standard dynamic programming arguments. For continuity of the value functions is important for applying Berge’s maximum theorem and measurable selection theorems [16, Proposition D5(a)] or [8, Corollary 2.3(iv)].
Let Let us take and consider a history Then Nature deals with a one-step MDP, and, in view of (A.1) and (A.6), the policy is optimal for the one-step problem if and only if for
Since the formulae
hold for every every policy every policy if and only if the policy satisfies condition (A.10) for So, if is an optimal policy for Nature, then (A.10) holds for In addition, according to [16, Proposition D5(a)] or its generalization [8, Corollary 2.3(iv)], there exists a Borel function such that for all and Thus, at the last step, an optimal policy for Nature can be chosen in the form of a one-step Markov policy described in the previous sentence. In addition, the function is continuous. Since each set is finite, in view of (A.2), the function is continuous.
If Nature plays an optimal policy , then for
and therefore the formulae
hold for all satisfying (A.10), all and all if and only if the policy satisfies the condition
Thus, for and for an arbitrary history a policy for the DM (policy for Nature), the policy (policy ) is optimal for the horizon problem if and only if condition (A.14) (condition (A.10)) holds.
Let us consider a nonnegative integer and make the following induction assumption. Assume that there are policies and such that, for each integer satisfying and for each history the policies and are optimal for the -horizon problem if and only if they satisfy conditions (A.14) and (A.10) respectively, and, in addition, if conditions (A.14) and (A.10) both hold for this then for all
Let and be the set of policies such that, for each integer satisfying , for each history conditions (A.14) and (A.10) hold respectively
The same proof as for imply that for and for a history inequality (A.11) holds for all policies and if and only if condition (A.10) holds. In addition, because of the same selection theorem, at time the policy can be selected Markov and nonrandomized. Similarly, (A.13) holds for a policy for all for all for all if and only if condition (A.14) holds. Thus, (i) – (iii) are proved by induction for .
Let Then a policy () is persistently optimal if and only if for every it is optimal for the -horizon problem with the terminal cost This is true because uniformly in In addition, and for all for this terminal cost. This implies (i) and (ii) for (iv) holds because for an infinite-horizon stationary policies are persistently optimal if and only if they are optimal. ∎
@@
Let us fix and let and be optimal Markov policies for the DM and Nature whose existence is stated in Theorem A.1. Let us fix As follows from the definition of optimal policies,
@@
The definition of persistently equilibrium pairs of policies for games with perfect information is consistent with the definition of persistently optimal policies for MDPs. An MDP can be viewed as a game with perfect information when the action space for Nature is a singleton, and the parameter is omitted from notations. A Markov policy is persistently optimal for an -horizon discounted MDP, or with the state space and with the expected total discounted costs if for all and for all
In particular, persistently optimal policies exist for an MDP if the state space is a Borel subset of a Polish space, the sets of all actions and the sets of all available actions A(x), are nonempty compact subsets of a Polish space, the set-valued mapping is upper semicontinuous, one step costs are bounded and continuous, and transition probabilities are weakly continuous in This fact follows from [9, Theorem 2], where a more general MDP is considered. In addition, the optimality equations for finite and infinite-step problems are
where is the given bounded lower semicontinuous function on Optimality equations define the sets of optimal actions
A Markov policy is persistently optimal for an -horizon problem with or if and only if for all nonnegative integers satisfying where, for an integer satisfying A deterministic policy is optimal for an infinite-horizon problem if and only if for all

### v4

@@ 3. The resulted finite sequence where and represents the concave piecewise linear function
Subroutine 1 is the major step in recursive computations of value functions for see the following section.
@@
This section describes the properties of the value functions and for DRMDP1 defined in equations (4.7), where or It also describes the properties of the sets of optimal actions for the DM. We recall that the function is concave and continuous in because the function is continuous in and because Assumption 5.1 holds. We observe that, if then the function is linear in
We observe that for fixed and or the first equation in (4.7) is a particular instance of equation (6.1) with with the variable in (6.1) substituted with with and with
@@
The state space is infinite, and functions and can be computed by discretizing the interval They can be also computed exactly if the state and action sets and the horizon length are not too large. For each computing can be done by Subroutine 1 described at the end of Section 6. According to the second equation in (4.7), for each functions are lower envelops in of piecewise-linear concave in functions and their computation is a standard task in computational geometry.
@@
Proof of Theorem 5.2.
For the given time horizon or and for a given sequence Algorithm DCVaR sequentially generates a finite or infinite sequence of optimal actions. Let be the states of the system and be the tail risk levels, and the values are not observed by the DM when In addition to the parameters of the model, the algorithm also uses the value functions , if or it uses the value function if
The parameter takes two values: 1 and 0. means that at the current time epoch the tail risk level is known, and indicates that it is not known. However, in the second case the algorithm finds an interval of possible risk levels, and this is sufficient to choose an optimal action.
At each iteration the algorithm chooses an action . In view of Theorem 4.5, in order to prove the theorem, it is sufficient to show that for all nonnegative integer for every persistently optimal policy played by Nature.
The proof is based on induction in According to step 1, Let an optimal action be selected at an epoch by step 1 or 2.4 of the algorithm. In view of (4.7),
Formula (6.6) implies
@@
If there are multiple points such that that is, the case takes place, then concavity of implies that there is a maximal interval such that the function is linear in and its slope is According to Lemma 7.2, for all the sets coincide, and for Therefore, step 2.1 can be skipped, and future calculations do not depend on the particular value So, in spite of the fact that the DM does not know the states when The algorithm either calculates values or can calculate intervals to which values belong, and optimal action sets coincides for tail risk levels, which are internal points these intervals. Actions from these set are also optimal for the extreme points and of this interval. ∎
@@ 9 Extension to Stochastic Cost Functions
In this section, we extend the results of this paper to random one-step cost functions with finite supports. Arbitrary random one-step costs can be approximated by such costs.
@@ Appendix A Results on Robust MDPs Used in this Paper
This appendix provides results on RMDPs used in this paper. In general, for robust optimization problems, some parameters of the model are unknown, and the goal is to select the best solutions for the worst possible values of unknown parameters. Starting from [20] and [27] there are numerous studies of RMDPs motivated by applications. For an RMDP, there is an additional parameter on which one-step costs and transition probabilities depend, and at each step Nature (sometimes called Player II) selects this parameter from a set of available parameters. The DM (sometines called Player I) chooses actions by knowing the past and current states and the past actions chosen by the DM and by Nature. Nature chooses actions by knowing this information and, in addition, by knowing the current action chosen by the DM. Such models were also studied under the names of two-person zero-sum games with perfect information (see e.g., [1, 12, 21]) and minimax control [14]. In RMDP papers [20, 27] the unknown parameter is the transition probability, and one-step costs depend only on states and actions selected by the DM. However, this specification, probably motivated by applications, does not simplify the theory needed in this paper. The models studied in [20, 27] are currently called the models satisfying the property of -rectangularity. Another property called -rectangularity was introduced in [44], and there are other rectangularity properties; see [13, 24, 15, 43]. Models with -rectangularity are relevant to two-person zero-sum stochastic games [21, 41], when the second player does not know the current action selected by the first player.
@@
An RMDP is defined by a tuple Here is the state space, and there are two players: the DM and Nature. There are two decision sets and where is the set of actions for the DM, and is the set of actions for Nature. We assume that , and are nonempty Borel subsets of Polish (complete, separable, metric) spaces. The sets where are assumed to be nonempty and finite, and are the sets of actions available to the DM at states For each pair where and a set of feasible actions of Nature is defined. We assume that the sets are nonempty and compact subsets of for all and for all We also assume that is a Borel subset of and the set-valued mapping is continuous. Continuity of implies Borel measurability of According to the Arsenin-Kunugui measurable selection theorem, each of the graphs and contains at least one Borel measurable function. This fact, which is needed for the existence of policies, also follows from the Kuratowski and Ryll-Nardzewski measurable selection theorem.
@@
where is a bounded continuous function (the terminal payoff), and for the infinite horizon
@@
where, in view of Berge’s maximum theorem, each function is continuous, and therefore is a continuous function. Equations (A.1) and (A.2) imply the minimax equation
Let Then all continuous functions are uniformly bounded by where Banach’s fixed point theorem applied to the space of uniformly bounded continuous functions on implies that the minimax operator applied in (A.3) to the function has a unique fixed point which is a bounded continuous function, and as In particular
@@
minimizing the right-hand side of (A.2), and for Nature, for and
maximizing the right-hand side of (A.1). The sets are finite. As follows from the Berge maximum theorem, the sets are compact.
For an -horizon problem with or a pair of policies is called an equilibrium if for all policies and
Theorem A.1.
For a given horizon or if two pairs of policies and are equilibria, then and are equilibria too, and for all and
Proof.
The definition of an equilibrium implies that for all This equality and the definition of equilibria imply the conclusions of the theorem. ∎
If a pair is an equilibrium for some then is called an optimal policy for the DM, and is called an optimal policy for Nature.
In order to formulate structural properties of optimal policies, we need to introduce additional notations and definitions. Let us fix a horizon or
Let the DM and Nature reset the clock to 0 at time Let the history prior to time be , and let and be the histories the DM and Nature observe respectively during the following units of time after the clock is reset to 0. If the clock is not not reset, then at time the history would be for the DM and for Nature.
@@
are the policies the DM and Nature will use respectively following their original policies and if the clock is reset to 0 at time
Definition A.2.
A policy for the DM ( for Nature) is called perfectly optimal for an -horizon problem, if it is optimal and, for each positive integer and for each the policy () is optimal for the horizon
Definition A.3.
A policy for Nature is called persistently optimal for an -horizon problem, if
for all nonnegative integers all and all
We have that a nonrandomized policy for Nature is persistently optimal if and only if for all nonnegative integers all and all A Markov policy for Nature is persistently optimal if and only if for all nonnegative integers all and all For a deterministic optimal policy for Nature is persistently optimal if and only if for all and all
@@
a policy for the DM is perfectly optimal if and only if for all nonnegative integers and for all
a persistently optimal policy for Nature is perfectly optimal;
there exists a Markov perfectly optimal policy for the DM, and a Markov policy for the DM is perfectly optimal if and only if for all nonnegative integers and for all
there exists a Markov persistently optimal policy for Nature, and is a Markov persistently optimal policy for Nature if and only if for all nonnegative integers all and all
if and are optimal policies for the DM and Nature respectively, then
for there exist deterministic optimal policies for the DM and Nature, and all deterministic optimal policy for the DM and for Nature are perfectly optimal;
for a deterministic policy for the DM is optimal if and only if for all and every deterministic persistently optimal policy for Nature is optimal.
@@
Let Let be a policy for the DM satisfying the conditions stated in and let be a persistently optimal policy for the Nature. Let us prove that the policies and are perfectly optimal for the DM and Nature respectively. For a history let us denote by its part that does not include the last state, that is, We also define for all policies. Straightforward backward calculations imply that if for and for all then for all In particular, (A.7) holds when If and then backward calculations imply that
Thus, and are perfectly optimal. So, statement and the sufficient condition in are proved. Since (A.7) holds for persistently optimal policies and considered above, Theorem A.1 implies that (A.7) holds for any pair of equilibrium policies. Thus (iii) is proved.
Let us prove the necessary condition in Let be a persistently optimal policy for the DM, and it does not satisfy the conditions stated in Let be a persistently optimal policy for Nature. In view of , be a perfectly optimal policy for Nature. Let us consider the largest nonnegative integer such that for some Then backward calculations imply that and thus the policy is not perfectly optimal. This contradiction implies the necessary condition in Statement follows from when is a Markov policy. Statement follows from definitions of persistently optimal policies and Markov policies for Nature.
Thus, statements are proved for Let us prove them for To prove let us consider a policy for the DM such that for all and all We also choose a persistently optimal policy for Nature for Let us consider a finite-horizon problem with the terminal value Then finite-horizon values for this problem with horizon are equal to and the sets and for this problem are equal to and respectively. Statements for a finite horizon implies that is perfectly optimal for the DM, and is perfectly optimal for Nature for all However, uniformly in , and this implies that policies and are perfectly optimal for the DM and Nature respectively for In particular, statement (iii) holds. To prove the necessary condition in let be a perfectly optimal policy for the DM. Then is perfectly optimal for every finite-horizon problem described in this paragraph. Thus, for all and all In view of Theorem A.1, statements and imply In particular, a deterministic policy for the DM )is perfectly optimal if and only if for all A deterministic policy for Nature is persistently optimal if and only if for all ∎
Corollary A.5.
For or if is an optimal policy for Nature, and is a policy for the DM satisfying the condition for all nonnegative integers then the following statements hold:
(ii) for all nonnegative integers and there exists a persistently optimal policy for Nature such that for all nonnegative integers and
Proof.
(i) Let be a perfectly optimal policy for the DM. Then, in view of Theorem A.4, for all and
Let us define the perfectly optimal policy for the DM: for nonnegative integer and for
Since for all nonnegative integer the probability measures and coincide. Therefore,
(ii) Let be a persistently optimal policy for Nature. Then, in view of Theorem A.4, for all
Let us define the persistently optimal policy for Nature: for nonnegative integer for and for
Let assume that it is possible that for some nonnegative integer Then
Thus, while the DM plays the perfectly optimal policy the optimal policy for Nature can be improved on the set of that has a positive probability if Nature switches to the policy Standard arguments imply that which is impossible since and are optimal policies. The last equality in (ii) follows from (i) since a persistently optimal policy is optimal. ∎
@@
Let us fix and let and be optimal policies for the DM and Nature whose existence is stated in Theorem A.4. Let us fix As follows from the definition of optimal policies,
@@
The following statement generalizes Theorem A.7.
Corollary A.8.
Let or and let and be subsets of policies for the DM and Nature such that there exist an optimal policy for the DM and an optimal policy for Nature. Then
Proof.
We observe that, since an optimal policy for
Therefore, similarly to the proof of Theorem A.7, for
and